OpenStudy (anonymous):

Find the sum of the first 15 terms of the arithmetic sequence whose first term is 6 and whose 15th term is 104.

OpenStudy (anonymous):

= (104+6)*15/2 = 110*15/2 = 55*15 = 825

OpenStudy (anonymous):

yes 825

OpenStudy (anonymous):

thank u!!!!!!